# DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation

### v1

@@ 5 Conclusion
In this work, we presented DAGLFNet, a pseudo-image-based network for point cloud semantic segmentation, designed to address challenges caused by point cloud sparsity and occlusions. By integrating local and global features, employing a multi-branch feature extraction mechanism, and leveraging deep feature-guided attention fusion, DAGLFNet achieves robust and accurate segmentation in long-range, sparse, and complex scenarios. Extensive experiments on benchmark datasets, including SemanticKITTI and nuScenes, demonstrate that the network attains competitive performance while maintaining real-time efficiency and enabling deployment on embedded devices.
We also identify a limitation of DAGLFNet in extremely sparse or occluded regions, where the model may misclassify semantically similar structures such as terrain and sidewalks, indicating that further improvement in leveraging fine-grained geometric and semantic context is needed in challenging environments.
Overall, DAGLFNet provides efficient and accurate segmentation for sparse and complex point clouds, balancing real-time performance and embedded deployment capability. Future research could focus on enhancing network understanding of geometric and semantic context in highly sparse and occluded areas, further improving robustness and accuracy in complex scenarios.

### v2

@@ 5 Conclusion
In this work, we introduce DAGLFNet, a pseudo-image-based framework for point cloud semantic segmentation, specifically engineered to mitigate challenges arising from point cloud sparsity and occlusions. By fusing local and global features, employing a multi-branch feature extraction mechanism, and utilizing deep feature-guided attention fusion, DAGLFNet achieves robust segmentation performance in long-range, sparse, and complex environments. Comprehensive evaluations on benchmark datasets—including SemanticKITTI and nuScenes—demonstrate that the network attains competitive accuracy while preserving real-time efficiency and suitability for edge deployment.
However, DAGLFNet exhibits limitations in regions of extreme sparsity or heavy occlusion, where it may misclassify semantically similar structures such as terrain and sidewalks. This underscores the need for enhanced modeling of fine-grained geometric and semantic context in challenging environments.
Overall, DAGLFNet delivers efficient and accurate segmentation for sparse and complex point clouds, striking a balance between real-time performance and embedded deployment capabilities. Future work should prioritize improving the network’s understanding of geometric and semantic context in highly sparse and occluded areas to further enhance robustness and accuracy in complex scenarios.